A ring is rotated about diametric axis in a uniform magnetic field perpendicular to the plane of the ring. If initially the plane of the ring is perpendicular to the magnetic field. Find the instant of time at which EMF will be maximum & minimum respectively :

Grade:12th pass

2 Answers

Arun
4 years ago
 ω = 2π/T = π/5 When ωt = π/2 ∵ ϕ will be minimum ∴ e will be maximum When ωt = π ∵ ϕ will have maximum ∵ e will be minimum
Vikas TU
4 years ago
 
 
ω = 2π/T = π/5 When ωt = π/2 ∵ ϕ will be minimum ∴ e will be maximum
t = 2.5 sec 
wt = pi 
Phi will be max, e will be minimum 
t = 5 sec